# Deep Neural Networks for Cross-Energy Particle Identification at RHIC and LHC

### v1

@@ 7 Conclusion
This paper demonstrates the ability of a fundamentally simple deep neural network trained exclusively on simulated LHC proton-proton collisions ($\sqrt{s}=13\,\mathrm{TeV}$), to generalize to RHIC data at a much lower center-of-mass energy. This generalization is achieved without any fine-tuning, weight adjustment or transfer learning. The model maintains an accuracy above 91% at all momentum cuts, even in the complex $p_{T}>7\,\mathrm{GeV}/c$ region. For the RHIC set under the same conditions, it achieves an accuracy of 96.23%, which is quite notable as the model was not trained on any $\sqrt{s}=200\,\mathrm{GeV}$ examples. The observed difference in the model’s performance is consistent with expectations, given the higher kinematic reach of the LHC. The ability of the model to generalize seamlessly to a lower center-of-mass energy highlights its robustness and suggests that it captures the underlying physics rather than overfitting to the kinematics of each regime.
These findings open the door to applying experimental simulation models in different energy domains beyond their original ones, especially in scenarios where training data is limited or simply unavailable. Future work will focus on introducing more detector-level features and employing more sophisticated network architectures, such as combined networks to handle the noisy nature of real experimental data.

### v3

@@ 7 Conclusion
This study demonstrates the ability of a deep neural network trained exclusively on simulated LHC proton-proton collisions ($\sqrt{s}=13\,\mathrm{TeV}$), to generalize effectively to RHIC data at a much lower center-of-mass energy. This generalization is achieved without any fine-tuning, weight adjustment or transfer learning. The model maintains an accuracy above 91% at all momentum cuts, even in the complex $p_{T}>7\,\mathrm{GeV}/c$ region. For the RHIC set under the same conditions, it achieves an accuracy of 96.23%, which is quite notable as the model was not trained on any $\sqrt{s}=200\,\mathrm{GeV}$ examples.
The observed difference in the model’s performance is consistent with expectations, given the higher kinematic reach of the LHC. Detailed analysis of per-class accuracy, particularly for kaons and electrons, reveals that the misidentification of particles is due to physical constraints related to production dynamics. The reduced classification accuracy for $K^{+}$ in high $p_{T}$ regions, especially at LHC energies, can be traced to the leading particle effect. This results in a broader and more energetic $p_{T}$ distribution relative to $K^{-}$, which in turn makes classification much more complex. Furthermore, complete misidentification of electrons at high $p_{T}$ arises from significant kinematic overlap with the much more common $\pi^{+}$ in the absence of discriminating detector-level features. The model’s consistent performance across different center-of-mass energies suggests that it captures the physically meaningful features, rather than overfitting to the specific kinematics of the training energy scale.
The model’s robust performance from LHC regime to RHIC test conditions is far from trivial even within a controlled and simulated environment due to differences in the kinematic distributions and production dynamics, as revealed by the confusion matrices. This study represents a first step and a proof of concept for assessing how deep learning models trained on higher center-of-mass energies behave when evaluated on lower collision energy data. The results warrant further investigation in real experimental settings to evaluate its practical viability. Applying this approach to real data naturally introduces multiple sources of uncertainty, including but not limited to detector resolution, reconstruction effects such as tracking inefficiencies, and domain shift between simulation and experiment. These challenges will need to be properly addressed in future studies.
A key direction for future application is to enhance the current kinematic feature set with crucial detector-level observables, to address the model’s difficulty in classifying certain particle types. On another front, future studies could explore more advanced network architectures which are better suited to deal with rich detector outputs. Prime candidates include convolutional neural networks which are ideal for energy deposition in calorimeters, ionization energy loss in TPC, and shower profiles. While CNNs represent a natural next step for analyzing spatial detector data, graph neural networks or transformer networks may offer more flexibility when dealing with complex topologies, sparse input features, and heterogeneous data types.